# Semantic Image Synthesis via Class-adaptive Cross-attention

### v2

@@ III Architecture
The proposed architecture (see Fig. 1) is called Class-Adaptive Cross-Attention, (CA)${}^{2}$-SIS, and is composed by three main modules: (a) a Multi-Resolution Style Encoder $\mathcal{E}_{s}$, with Grouped Convolutions, Group Normalization layers and skip connections, that is used to extract style features from an RGB reference, (b) a Mask Embedder $\mathcal{E}_{m}$ that extracts a latent representation separately for each semantic class and, finally, (c) a Cross-Attention Generator $\mathcal{G}$ that exploits the attention mechanism to inject the multi-resolution style codes inside the network in order to control the generated image. Additionally, a Discriminator $\mathcal{D}$ is employed during training to enforce the adversarial loss.
Figure 1: (CA)${}^{2}$-SIS architecture: style codes are extracted using a Multi-Resolution Style Encoder $\mathcal{E}_{s}$ equipped with grouped convolutions; the Mask Embedder $\mathcal{E}_{m}$ embeds each of the semantic mask parts into a set of latent codes; finally, these codes are fed to the Cross-Attention Generator $\mathcal{G}$ that is conditioned with the style codes thanks to the cross-attention mechanism. Additionally, the semantic mask is also used in the cross-attention layer to calculate the attention loss $\mathcal{L}_{att}$ in order to push each attention map to follow the mask shape.
@@ IV-A Reconstruction
Figure 5: Comparison between V-INADE, SEAN and (CA)${}^{2}$-SIS when transferring the style of all face parts. Our method convincingly generates realistic results even if a specific style i.e. opposite ear in the top row, eyeglasses in the middle row, and teeth in the bottom row, are absent in the reference image.
@@ IV-B Style transfer
Figure 6: Style transfer of face parts. (CA)${}^{2}$-SIS can successfully perform local style manipulation even if no spatial information is used in the generator.
@@ IV-D Comparison with StyleGAN-based methods
Using GAN-Inversion, remarkable results have been achieved. Yet, a question is whether the inversion process is sufficiently accurate to (i) produce an image where the identity of the subject is sufficiently preserved, (ii) maintain the level of quality and realism achieved in a purely generative setting and (iii) be applied to out-of-domain data i.e. faces not included in CelebA-HQ. In this section, we aim at comparing the results of our method against the most recent alternative based on StyleGAN, which is SemanticStyleGAN [30]. This specific approach also performs local face editing based on a semantic mask. Results are reported in Table II. The performance of SemanticStyleGAN drops dramatically when applied in a GAN-inversion setting. The quality of the generated images decreases significantly as evidenced by the FID score which increases to 26.83 from 6.42 (value taken from [30]) of the fully generative setting. Also, the ability of retaining the identity of the subject is compromised. We measure the latter using the Face Recognition Similarity (FRS) score: it is computed by embedding both real and reconstructed (inverted) images using an InceptionV3 pre-trained on VGG-Face2 dataset [31], and then computing the cosine similarity between embedding pairs. Our (CA)${}^{2}$-SIS scores a way higher similarity, indicating the identity is more effectively preserved. This is an essential feature for methods that aim at accurate editing of real images.
Methods based on StyleGAN are extremely promising in the generation task, yet these results evidence that they still fall short when manipulating a real image is the goal as they fail to apply the inversion process in a sufficiently accurate way. Another drawback is related to the worse generalization to out-of-distribution data. In Figure 12, we qualitatively compare (CA)${}^{2}$-SIS, SemanticStyleGAN and SEAN on some images from FFHQ [22], a commonly used dataset for training generative models. Semantic masks are not provided in FFHQ, so we employed the same face parser of Section IV to generate them. All the three models were trained on CelebA-HQ. SemanticStyleGAN fails dramatically in maintaining the identity traits of the subjects, mostly for severely under-represented classes such as the eyeglasses (top row), babies (middle row) or non-frontal faces (bottom row). SEAN instead can better preserve the identity and handle unseen samples, but results less robust to noise in the semantic masks. Our model more robustly handles detection noise as compared to SEAN. This is another advantageous result of embedding the mask into a latent vector and relying on attention mechanism.
@@ IV-E Subjective Evaluation
Figure 13: Subjective evaluation. Both charts show (CA)${}^{2}$-SIS is preferred by users, who selected our results as the most realistic in the majority of cases.
@@ IV-F Ablation study
Style Encoder FID $\downarrow$ Ours - SEAN $\mathcal{E}_{s}$ 17.81 Ours w/o GC 16.42 (CA)${}^{2}$-SIS 15.84 Generator FID $\downarrow$ Ours w/o $\mathcal{E}_{m}$ 18.65 Ours w/o CA 18.14 (CA)${}^{2}$-SIS 15.84
@@
Method FID $\downarrow$ FID-Sh $\downarrow$ SEAN 18.72 20.04 V-INADE 17.49 31.16 (CA)${}^{2}$-SIS 15.84 15.90
@@ IV-G Diversity-driven generation
The main goal of this work was to explore a new, alternative architecture for semantic image synthesis guided by a reference style image. However, the proposed generator architecture based on spatial transformer layers in place of SPADE-like normalization is general in its purpose and can be potentially used also in a fully-generative setting. Clearly, the latter requires specific training strategies to force both quality and diversity in the generated images, which is beyond the scope of this work. Nevertheless, in this section we report the results of an additional set of experiments aimed at showing that we can train (CA)${}^{2}$-SIS in the diversity-driven setting and obtain results that are on-par with respect to state-of-the-art methods specifically tailored for such task.
@@
Results on CelebMask-HQ are shown in Table V, in comparison with three state-of-the-art approaches, namely INADE [16], SC-GAN [3] and SemanticDiffusion [8]. Note that, differently from the methods in Table I, these are purely diversity-driven and do not use any reference images. In terms of image quality (FID), our method’s performance are comparable to both previous GANs equipped with SPADE layers (INADE), other custom modules such as conditional convolutions (SC-GAN) and SemanticDiffusion, which also employs SPADE to condition the diffusion process with the semantic mask. In terms of diversity (LPIPS [33]) instead, SemanticDiffusion stands out. About the latter, metrics were computed on images generated with $1,000$ diffusion steps to ensure the best possible quality. However, generating the whole test set (2K images) in this way, took approximately 11 days on a NVIDIA A100 GPU (8 minutes per image). We also tried using the DDIM [34] technique to reduce the diffusion steps. Using only 50 steps, generating a single image takes approximately 10 seconds, yet the FID grows from $19.22$ to $32.96$. Our (CA)${}^{2}$-SIS instead takes $0.07$ seconds to generate one image. Whereas using a diffusion models provided excellent results, the computational burden makes its usage quite unpractical. Moreover, our solution allows for a nice feature not shown by previous methods, that is generating diverse and disentangled styles for different semantic classes. For example, one can keep the style of some parts fixed while only generating that of a specific class, such as skin or hair, allowing for a detailed generation control over local classes. Some qualitative results are shown in Figure 15.

### v3

@@ III Architecture
The proposed architecture (see Fig. 1) is called Class-Adaptive Cross-Attention, (CA)2-SIS, and is composed by three main modules: (a) a Multi-Resolution Style Encoder $\mathcal{E}_{s}$, with Grouped Convolutions, Group Normalization layers and skip connections, that is used to extract style features from an RGB reference, (b) a Mask Embedder $\mathcal{E}_{m}$ that extracts a latent representation separately for each semantic class and, finally, (c) a Cross-Attention Generator $\mathcal{G}$ that exploits the attention mechanism to inject the multi-resolution style codes inside the network in order to control the generated image. Additionally, a Discriminator $\mathcal{D}$ is employed during training to enforce the adversarial loss.
Figure 1: (CA)2-SIS architecture: style codes are extracted using a Multi-Resolution Style Encoder $\mathcal{E}_{s}$ equipped with grouped convolutions; the Mask Embedder $\mathcal{E}_{m}$ embeds each of the semantic mask parts into a set of latent codes; finally, these codes are fed to the Cross-Attention Generator $\mathcal{G}$ that is conditioned with the style codes thanks to the cross-attention mechanism. Additionally, the semantic mask is also used in the cross-attention layer to calculate the attention loss $\mathcal{L}_{att}$ in order to push each attention map to follow the mask shape.
@@ IV-A Reconstruction
Figure 5: Comparison between V-INADE, SEAN and (CA)2-SIS when transferring the style of all face parts. Our method convincingly generates realistic results even if a specific style i.e. opposite ear in the top row, eyeglasses in the middle row, and teeth in the bottom row, are absent in the reference image.
@@ IV-B Style transfer
Figure 6: Style transfer of face parts. (CA)2-SIS can successfully perform local style manipulation even if no spatial information is used in the generator.
@@ IV-D Comparison with StyleGAN-based methods
Using GAN-Inversion, remarkable results have been achieved. Yet, a question is whether the inversion process is sufficiently accurate to (i) produce an image where the identity of the subject is sufficiently preserved, (ii) maintain the level of quality and realism achieved in a purely generative setting and (iii) be applied to out-of-domain data i.e. faces not included in CelebA-HQ. In this section, we aim at comparing the results of our method against the most recent alternative based on StyleGAN, which is SemanticStyleGAN [30]. This specific approach also performs local face editing based on a semantic mask. Results are reported in Table II. The performance of SemanticStyleGAN drops dramatically when applied in a GAN-inversion setting. The quality of the generated images decreases significantly as evidenced by the FID score which increases to 26.83 from 6.42 (value taken from [30]) of the fully generative setting. Also, the ability of retaining the identity of the subject is compromised. We measure the latter using the Face Recognition Similarity (FRS) score: it is computed by embedding both real and reconstructed (inverted) images using an InceptionV3 pre-trained on VGG-Face2 dataset [31], and then computing the cosine similarity between embedding pairs. Our (CA)2-SIS scores a way higher similarity, indicating the identity is more effectively preserved. This is an essential feature for methods that aim at accurate editing of real images.
Methods based on StyleGAN are extremely promising in the generation task, yet these results evidence that they still fall short when manipulating a real image is the goal as they fail to apply the inversion process in a sufficiently accurate way. Another drawback is related to the worse generalization to out-of-distribution data. In Figure 12, we qualitatively compare (CA)2-SIS, SemanticStyleGAN and SEAN on some images from FFHQ [22], a commonly used dataset for training generative models. Semantic masks are not provided in FFHQ, so we employed the same face parser of Section IV to generate them. All the three models were trained on CelebA-HQ. SemanticStyleGAN fails dramatically in maintaining the identity traits of the subjects, mostly for severely under-represented classes such as the eyeglasses (top row), babies (middle row) or non-frontal faces (bottom row). SEAN instead can better preserve the identity and handle unseen samples, but results less robust to noise in the semantic masks. Our model more robustly handles detection noise as compared to SEAN. This is another advantageous result of embedding the mask into a latent vector and relying on attention mechanism.
@@ IV-E Subjective Evaluation
Figure 13: Subjective evaluation. Both charts show (CA)2-SIS is preferred by users, who selected our results as the most realistic in the majority of cases.
@@ IV-F Ablation study
Style Encoder FID $\downarrow$ Ours - SEAN $\mathcal{E}_{s}$ 17.81 Ours w/o GC 16.42 (CA)2-SIS 15.84 Generator FID $\downarrow$ Ours w/o $\mathcal{E}_{m}$ 18.65 Ours w/o CA 18.14 (CA)2-SIS 15.84
@@
Method FID $\downarrow$ FID-Sh $\downarrow$ SEAN 18.72 20.04 V-INADE 17.49 31.16 (CA)2-SIS 15.84 15.90
@@ IV-G Diversity-driven generation
The main goal of this work was to explore a new, alternative architecture for semantic image synthesis guided by a reference style image. However, the proposed generator architecture based on spatial transformer layers in place of SPADE-like normalization is general in its purpose and can be potentially used also in a fully-generative setting. Clearly, the latter requires specific training strategies to force both quality and diversity in the generated images, which is beyond the scope of this work. Nevertheless, in this section we report the results of an additional set of experiments aimed at showing that we can train (CA)2-SIS in the diversity-driven setting and obtain results that are on-par with respect to state-of-the-art methods specifically tailored for such task.
@@
Results on CelebMask-HQ are shown in Table V, in comparison with three state-of-the-art approaches, namely INADE [16], SC-GAN [3] and SemanticDiffusion [8]. Note that, differently from the methods in Table I, these are purely diversity-driven and do not use any reference images. In terms of image quality (FID), our method’s performance are comparable to both previous GANs equipped with SPADE layers (INADE), other custom modules such as conditional convolutions (SC-GAN) and SemanticDiffusion, which also employs SPADE to condition the diffusion process with the semantic mask. In terms of diversity (LPIPS [33]) instead, SemanticDiffusion stands out. About the latter, metrics were computed on images generated with $1,000$ diffusion steps to ensure the best possible quality. However, generating the whole test set (2K images) in this way, took approximately 11 days on a NVIDIA A100 GPU (8 minutes per image). We also tried using the DDIM [34] technique to reduce the diffusion steps. Using only 50 steps, generating a single image takes approximately 10 seconds, yet the FID grows from $19.22$ to $32.96$. Our (CA)2-SIS instead takes $0.07$ seconds to generate one image. Whereas using a diffusion models provided excellent results, the computational burden makes its usage quite unpractical. Moreover, our solution allows for a nice feature not shown by previous methods, that is generating diverse and disentangled styles for different semantic classes. For example, one can keep the style of some parts fixed while only generating that of a specific class, such as skin or hair, allowing for a detailed generation control over local classes. Some qualitative results are shown in Figure 15.